# PaSa: An LLM Agent for Comprehensive Academic Paper Search

### v1

@@ Abstract
We introduce PaSa, an advanced Paper Search agent powered by large language models. PaSa can autonomously make a series of decisions, including invoking search tools, reading papers, and selecting relevant references, to ultimately obtain comprehensive and accurate results for complex scholarly queries. We optimize PaSa using reinforcement learning with a synthetic dataset, AutoScholarQuery, which includes 35k fine-grained academic queries and corresponding papers sourced from top-tier AI conference publications. Additionally, we develop RealScholarQuery, a benchmark collecting real-world academic queries to assess PaSa performance in more realistic scenarios. Despite being trained on synthetic data, PaSa significantly outperforms existing baselines on RealScholarQuery, including Google, Google Scholar, Google with GPT-4 for paraphrased queries, chatGPT (search-enabled GPT-4o), GPT-o1, and PaSa-GPT-4o (PaSa implemented by prompting GPT-4o). Notably, PaSa-7B surpasses the best Google-based baseline, Google with GPT-4o, by 37.78% in recall@20 and 39.90% in recall@50. It also exceeds PaSa-GPT-4o by 30.36% in recall and 4.25% in precision. Model, datasets, and code are available at https://github.com/bytedance/pasa.
@@
Yichen He1 Guanhua Huang1  Peiyuan Feng1  Yuan Lin1 Yuchen Zhang1Hang Li1Weinan E2 1ByteDance Research  2Peking University {hyc,huangguanhua,fpy,linyuan.0}@bytedance.com, {zhangyuchen.zyc,lihang.lh}@bytedance.com, weinan@math.pku.edu.cn Demo: https://pasa-agent.ai
\$*\$\$*\$footnotetext: Equal contribution.$\dagger$\$$\dagger$\$footnotetext: Corresponding author.
@@ 1 Introduction
Although AutoScholarQuery only provides query and paper answers, without demonstrating the path by which scientists collect the papers, we can utilize them to perform RL training to improve PaSa. In addition, we design a new session-level PPO (Proximal Policy Optimization Schulman et al. (2017)) training method to address the unique challenges of the paper search task: 1) sparse reward: The papers in AutoScholarQuery are collected via citations, making it a smaller subset of the actual qualified paper set. 2) long trajectories: The complete trajectory of the Crawler may involve hundreds of papers, which is too long to directly input into the LLM context.
@@ 3.1 AutoScholarQuery
To construct AutoScholarQuery, we began by collecting all papers published at ICLR 2023, ICML 2023, NeurIPS 2023, ACL 2024, and CVPR 2024. For the Related Work section of each paper, we prompted GPT-4o Hurst et al. (2024) to generate scholarly queries, where the answers to these queries correspond to the references cited in the Related Work section. The prompt used is shown in Appendix E.1. For each query, we retained only the papers that could be retrieved on arXiv [^6] https://arxiv.org/, using their arxiv_id as the unique article identifier in the dataset. We adopt the publication date of the source paper as the query date. During both training and testing, we only considered papers published prior to the query date.
@@
Query: Could you provide me some studies that proposed hierarchical neural models to capture spatiotemporal features in sign videos? Query Date: 2023-05-02 Answer Papers: [1] TSPNet: Hierarchical Feature Learning via Temporal Semantic Pyramid for Sign Language Translation [2] Sign Language Translation with Hierarchical Spatio-Temporal Graph Neural Network
@@
Query: Which studies have focused on nonstationary RL using value-based methods, specifically Upper Confidence Bound (UCB) based algorithms? Query Date: 2023-08-10 Answer Papers: [1] Reinforcement Learning for Non-Stationary Markov Decision Processes: The Blessing of (More) Optimism [2] Efficient Learning in Non-Stationary Linear Markov Decision Processes [3] Nonstationary Reinforcement Learning with Linear Function Approximation Source: Provably Efficient Algorithm for Nonstationary Low-Rank MDPs, NeurIPS 2023
Query: Which studies have been conducted in long-form text generation, specifically in story generation? Query Date: 2024-01-26 Answer Papers: [1] Strategies for Structuring Story Generation [2] MEGATRON-CNTRL: Controllable Story Generation with External Knowledge Using Large-Scale Language Models Source: ProxyQA: An Alternative Framework for Evaluating Long-Form Text Generation with Large Language Models, ACL 2024
@@
Table 2: Statistics of AutoScholarQuery. $|P|$ and $|Q|$ represent the total number of papers and queries collected for each conference. $Ans(/Q)$ denotes the average number of answer papers per query. $Ans$-$50$ and $Ans$-$90$ refers to the 50th and 90th percentiles of answer paper counts per query.
To evaluate the quality of AutoScholarQuery, we sampled 100 query-paper pairs and assessed the rationality and relevance of each query and the corresponding paper. A qualified query should be meaningful and unambiguous. A qualified paper should match the requirements of the scholarly query. The author manually reviewed each pair, determining that 94.0% of the queries were qualified. Among these qualified queries, 93.7% had corresponding papers that were deemed relevant and appropriate.
@@ 3.2 RealScholarQuery
To evaluate PaSa in more realistic scenarios, we constructed RealScholarQuery, a test dataset consisting of 50 real-world research queries. After launching the demo of PaSa, we invited several AI researchers to use the system. From the queries they provided, we randomly sampled a subset of queries and manually filtered out overly broad topics (e.g., "multi-modal large language models," "video generation"). Ultimately, we collected 50 fine-grained and realistic queries.
For each query, we first manually gathered relevant papers. Subsequently, we used multiple methods to retrieve additional papers, including PaSa, Google, Google Scholar, ChatGPT (search-enabled GPT-4o), and Google paired with GPT-4o for paraphrased queries. The results from these methods were aggregated into a pool of candidate papers. Finally, professional annotators reviewed all candidate papers for each query, selecting those that met the specific requirements of the query to create the final set of relevant papers. The query date of all instances in RealScholarQuery is 2024-10-01. Table 12 in Appendix D provides examples from RealScholarQuery.
The annotators included professors from the Department of Computer Science at a top-tier university in China. On average, each query required the annotators to review 76 candidate papers. Given the high cost of the annotations, we completed this process for only 50 instances.
@@ 4.1 Overview
As illustrated in Figure 1, the PaSa system consists of two LLM agents: Crawler and Selector. The crawler reads the user’s query, generates multiple search queries, and retrieves relevant papers. The retrieved papers are added to a paper queue. The Crawler further processes each paper in the paper queue to identify key citations worth exploring further, appending any newly relevant papers to the paper list. The selector conducts a thorough review of each paper in the paper list to assess whether it fulfills the user’s query requirements.
@@
Figure 2: An example of the PaSa workflow. The Crawler runs multiple [Search] using diverse and complementary queries. In addition, the Crawler can evaluate the long-term value of its actions. Notably, it discovers many relevant papers as it explores deeper on the citation network, even when intermediate papers along the path do not align with the user query.
@@ 4.2 Crawler
For example, as Figure 2 shows, the agent begins by receiving a user query, incorporating it into its context, and initiating actions. If the token generated is [Search], the LLM continues to generate a search query, and the agent invokes a search tool to retrieve papers, which are then added to the paper list. If the token is [Expand], the LLM continues to extract a subsection name from the current paper in its context. The agent subsequently uses a parsing tool to extract all referenced papers within that subsection, adding them to the paper list. If the token is [Stop], the agent resets its context to the user query and information of the next paper in the paper queue. This information includes the title, abstract, and an outline of all sections and subsections.
The training process for the Crawler comprises two stages. In the first stage, we generate trajectories for a small subset of the training data and then perform imitation learning (see Appendix A.1 for details). In the second stage, reinforcement learning is applied. The details of the RL training implementation are described below.
@@
[…]
@@
A key challenge in training the Crawler with RL is the significant time required to sample a complete trajectory for a given query. This is due to each [Search] or [Expand] action adding multiple papers to the paper list, resulting in hundreds or even thousands of papers in the final paper queue.
To address this issue, we define a session as a sub-trajectory that begins with a session’s initial state and ends with the [Stop] action. We identify two types of session initial states: $S_{q}$, which includes only a query, and $S_{q+p}$, which consists of both a query and a paper.
Formally, we model the Crawler as a policy $\pi_{\theta}(a_{t}|s_{t})$. We partition the entire trajectory $\tau$ into a sequence of sessions: $(\tau_{t_{1}:t_{2}-1},\tau_{t_{2}:t_{3}-1},\cdots)$. Each session is $\tau_{t_{i}:t_{i+1}-1}=(s_{t_{i}},a_{t_{i}},\cdots,s_{t_{i+1}-1},a_{t_{i+1}-1})$, where the initial state $s_{t_{i}}$ is either belonging to type $S_{q}$ or $S_{q+p}$, and the final action $a_{t_{i+1}-1}$ is [STOP].
@@
$\hat{R}_{t}$ $=$ $\sum_{k=0}^{t_{i+1}-1-t}\gamma_{0}^{k}\bigg{[}r(s_{t+k},a_{t+k})$
$+\gamma_{1}\sum_{j=1}^{n_{t+k}}\hat{V}_{\phi}(S_{q+p_{j}})-\beta\cdot\log\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\rm sft}(a_{t}|s_{t})}\bigg{]}$
Here, $\gamma_{0}$ is the in-session discount factor, and $\gamma_{1}$ is the across-session discount factor. $\hat{V}_{\phi}(\cdot)$ is the value function model to approximate the state value. After executing $a_{t+k}$, the paper queue is updated to include the newly found papers $(p_{1},p_{2},\cdots,p_{n_{t+k}})$. Since the Crawler will subsequently initiate new sessions to process these additional papers, their associated reward-to-go should be incorporated into the return estimate. In addition, we include a per-token KL penalty term from the learned policy $\pi_{\theta}$ to the initial policy $\pi_{\rm sft}$ obtained through imitation learning at each token to mitigate over-optimization. This term is scaled by the coefficient $\beta$.
@@ 4.3 Selector
We perform imitation learning to optimize the Selector. See Appendix B for training data collection and training details.
@@ 5.1 Experimental Setting
The Selector was fine-tuned using the training dataset described in Appendix B. We conducted supervised fine-tuning for one epoch with a learning rate of 1e-5 and a batch size of 4. The training runs on 8 NVIDIA-H100 GPUs.
@@
The training process involves two stages. First, we perform imitation learning for 1 epoch on 12,989 training data with a learning rate of 1e-5 and batch size of 4 per device, using 8 NVIDIA H100 GPUs. In the second stage, we apply PPO training. To ensure stability, we first freeze the policy model and train the value model, followed by co-training both the policy and value models. The hyperparameters used during the training process are listed in the Table 4.
Name Value
$\alpha$ (Equation 1) 1.5
[…]
$\gamma_{0}$ (Equation 4.2) 1.0
$\gamma_{1}$ (Equation 4.2) 0.1
$\beta$ (Equation 4.2) 0.1
$\epsilon$ (Equation 5, Equation 4.2) 0.2
$\eta$ (Equation 8) 10
learning rate 1e-6
epoch per step 2
forward batch size 1
accumulate batch size 16
NVIDIA H100 GPU 16
policy freezing step 50
total step 250
Table 4: The hyperparameters used in PPO training.
During imitation learning, the model encounters 5,000 queries, while during the RL training phase, the model processes a total of 16,000 queries. For more details please refer to Appendix A.1 for the imitation learning data construction and Appendix A.2 for the PPO training data sampling.
@@
This function utilizes the LLM to predict a query based on the context, and then calls Google [^7] Accessed via the Google Search API provided by https://serper.dev. with the parameters site:arxiv.org and before:query_date, restricting search results by source and publication time.
@@
Table 5: Results on AutoScholarQuery test set.
@@ 5.2 Baselines and Evaluation
  • ChatGPT. We submit the scholar query to ChatGPT [^9] https://chatgpt.com, powered by search-enabled GPT-4o. Due to the need for manual query submission, we evaluate only 100 randomly sampled instances from the AutoScholarQuery test set.
  • GPT-o1. Prompt GPT-o1 to process the scholar query.
  • PaSa-GPT-4o. Prompt GPT-4o within the PaSa framework. It can perform multiple searches, paper reading, and citation network crawling.
We carefully designed prompts for all baselines and they are shown in Appendix E.1.
As shown in Figure 2, the crawling process of PaSa can be visualized as a paper tree. In practice, considering the computational expense, we limit the Crawler’s exploration depth (starting from the user query) to three for both PaSa-7b and PaSa-GPT-4o.
For Google-based baselines, we evaluate recall using Recall@20, Recall@50, and Recall@100 metrics for the top-20, top-50, and top-100 search results, respectively. For other baselines, we assess precision and recall for the final retrieved papers. Additionally, we compare the crawler’s recall between PaSa-GPT-4o and PaSa-7b.
5.3 Main results
We observe that using multiple ensembles of Crawler during inference can improve performance. Specifically, running Crawler twice during inference increased the Crawler recall by 3.34% on AutoScholarQuery, leading to the final recall improvement by 1.51%, with precision remaining similar.
As both the final decision-maker and auxiliary reward model in RL training for the Crawler, the performance of the Selector is crucial. To evaluate its effectiveness, we collected a dataset of 200 query-paper pairs, annotating whether each paper meets the query’s requirements. This dataset serves as the benchmark for evaluating the Selector (see Appendix C for details). We then compared our Selector against GPT-4o Hurst et al. (2024) and Qwen-2.5-7b Yang et al. (2024), as shown in Table 7. The results show that our Selector achieves an F1 score of 85%, outperforming GPT-4o by 5% and Qwen-2.5-7b by 30%. Additionally, when compared to a setting where reasoning precedes decision token generation, the performance is comparable. Lastly, the Selector’s precision reaches 95%, confirming its effectiveness as an auxiliary reward model for the Crawler RL training.

### v2

@@ Abstract
We introduce PaSa, an advanced Paper Search agent powered by large language models. PaSa can autonomously make a series of decisions, including invoking search tools, reading papers, and selecting relevant references, to ultimately obtain comprehensive and accurate results for complex scholar queries. We optimize PaSa using reinforcement learning with a synthetic dataset, AutoScholarQuery, which includes 35k fine-grained academic queries and corresponding papers sourced from top-tier AI conference publications. Additionally, we develop RealScholarQuery, a benchmark collecting real-world academic queries to assess PaSa performance in more realistic scenarios. Despite being trained on synthetic data, PaSa significantly outperforms existing baselines on RealScholarQuery, including Google, Google Scholar, Google with GPT-4o for paraphrased queries, ChatGPT (search-enabled GPT-4o), GPT-o1, and PaSa-GPT-4o (PaSa implemented by prompting GPT-4o). Notably, PaSa-7B surpasses the best Google-based baseline, Google with GPT-4o, by 37.78% in recall@20 and 39.90% in recall@50, and exceeds PaSa-GPT-4o by 30.36% in recall and 4.25% in precision. Model, datasets, and code are available at https://github.com/bytedance/pasa.
@@
Yichen He1  Guanhua Huang1   Peiyuan Feng1   Yuan Lin1 Yuchen Zhang1Hang Li1Weinan E2 1ByteDance Seed  2Peking University {hyc,huangguanhua,fpy,linyuan.0}@bytedance.com, {zhangyuchen.zyc,lihang.lh}@bytedance.com, weinan@math.pku.edu.cn Demo: https://pasa-agent.ai
[…]
@@ 1 Introduction
Although AutoScholarQuery only provides query and paper answers, without demonstrating the path by which scientists collect the papers, we can utilize it to perform RL training to improve PaSa. In addition, we design a new session-level PPO (Proximal Policy Optimization Schulman et al. (2017)) training method to address the unique challenges of the paper search task: 1) sparse reward: The papers in AutoScholarQuery are collected via citations, making it a smaller subset of the actual qualified paper set. 2) long trajectories: The complete trajectory of the Crawler may involve hundreds of papers, which is too long to directly input into the LLM context.
@@ 3.1 AutoScholarQuery
To construct AutoScholarQuery, we began by collecting all papers published at ICLR 2023, ICML 2023, NeurIPS 2023, ACL 2024, and CVPR 2024. For the Related Work section of each paper, we prompted GPT-4o Hurst et al. (2024) to generate scholarly queries, where the answers to these queries correspond to the references cited in the Related Work section. The prompt used is shown in Appendix H.1. For each query, we retained only the papers that could be retrieved on arXiv [^6] https://arxiv.org/, using their arxiv_id as the unique article identifier in the dataset. We adopt the publication date of the source paper as the query date. During both training and testing, we only considered papers published prior to the query date.
@@
To evaluate the quality of AutoScholarQuery, we sampled 100 query-paper pairs and assessed the rationality and relevance of each query and the corresponding paper. A qualified query should be meaningful and unambiguous. A qualified paper should match the requirements of the scholarly query. Detailed evaluation criteria are provided in Appendix. A. Three authors manually reviewed each pair, determining that 94.0% of the queries were qualified. Among these qualified queries, 93.7% had corresponding papers that were deemed relevant and appropriate.
Query: Could you provide me some studies that proposed hierarchical neural models to capture spatiotemporal features in sign videos? Query Date: 2023-05-02 Answer Papers: [1] TSPNet: Hierarchical Feature Learning via Temporal Semantic Pyramid for Sign Language Translation (2010.05468) [2] Sign Language Translation with Hierarchical Spatio-Temporal Graph Neural Network (2111.07258)
@@
Query: Which studies have focused on nonstationary RL using value-based methods, specifically Upper Confidence Bound (UCB) based algorithms? Query Date: 2023-08-10 Answer Papers: [1] Reinforcement Learning for Non-Stationary Markov Decision Processes: The Blessing of (More) Optimism (2006.14389) [2] Efficient Learning in Non-Stationary Linear Markov Decision Processes (2010.12870) [3] Nonstationary Reinforcement Learning with Linear Function Approximation (2010.04244) Source: Provably Efficient Algorithm for Nonstationary Low-Rank MDPs, NeurIPS 2023
Query: Which studies have been conducted in long-form text generation, specifically in story generation? Query Date: 2024-01-26 Answer Papers: [1] Strategies for Structuring Story Generation (1902.01109) [2] MEGATRON-CNTRL: Controllable Story Generation with External Knowledge Using Large-Scale Language Models (2010.00840) Source: ProxyQA: An Alternative Framework for Evaluating Long-Form Text Generation with Large Language Models, ACL 2024
@@
Table 2: Statistics of AutoScholarQuery. $|P|$ and $|Q|$ represent the total number of papers and queries collected for each conference. $Ans(/Q)$ denotes the average number of answer papers per query. AnsAns-5050 and AnsAns-9090 refers to the 50th and 90th percentiles of answer paper counts per query.
Figure 2: An example of the PaSa workflow. The Crawler runs multiple [Search] using diverse and complementary queries. In addition, the Crawler can evaluate the long-term value of its actions. Notably, it discovers many relevant papers as it explores deeper in the citation network, even when intermediate papers along the path do not align with the user query.
@@ 3.2 RealScholarQuery
To evaluate PaSa in more realistic scenarios, we constructed RealScholarQuery, a test dataset consisting of 50 real-world research queries. After launching the demo of PaSa, we invited several AI researchers to use the system. From the queries they provided, we randomly sampled a subset of queries and manually filtered out overly broad topics (e.g., "multimodal large language models," "video generation"). Ultimately, we collected 50 fine-grained and realistic queries.
For each query, we first manually gathered relevant papers to the best of our ability. To ensure comprehensive coverage, we then applied multiple methods to retrieve additional papers, including PaSa, Google, Google Scholar, ChatGPT (search-enabled GPT-4o), and Google paired with GPT-4o for paraphrased queries. As these methods also serve as baselines for comparison with PaSa, implementation details are deferred to Section 5.2. The results from all methods were aggregated into a pool of candidate papers. Finally, professional annotators reviewed all candidate papers for each query, selecting those that met the specific requirements of the query to create the final set of relevant papers. Annotation guidelines and quality control procedures are detailed in Appendix. B. The query date of all instances in RealScholarQuery is 2024-10-01. Table 9 in Appendix C provides an example from RealScholarQuery.
The annotators included professors from the Department of Computer Science at a top-tier university in China. On average, each query required the annotators to review 76 candidate papers. We paid \$4 per data entry (a query-paper pair), resulting in an average of \$304 per query. Given the high annotation cost, we completed this process for only 50 instances. On average, each query is associated with 15.82 answer papers. The 50th percentile of answer counts per query is 9, while the 90th percentile reaches 37.
@@ 4.1 Overview
As illustrated in Figure 1, the PaSa system consists of two LLM agents: Crawler and Selector. The crawler reads the user’s query, generates multiple search queries, and retrieves relevant papers. The retrieved papers are added to a paper queue. The Crawler further processes each paper in the paper queue to identify key citations worth exploring further, appending any newly relevant papers to the paper queue. The selector conducts a thorough review of each paper in the paper queue to assess whether it fulfills the user’s query requirements.
@@ 4.2 Crawler
For example, as Figure 2 shows, the agent begins by receiving a user query, incorporating it into its context, and initiating actions. If the token generated is [Search], the LLM continues to generate a search query, and the agent invokes a search tool to retrieve papers, which are then added to the paper queue. If the token is [Expand], the LLM continues to extract a subsection name from the current paper in its context. The agent then extracts all referenced papers within that subsection, adding them to the paper queue. If the token is [Stop], the agent resets its context to the user query and information of the next paper in the paper queue. This information includes the title, abstract, and an outline of all sections and subsections.
The training process for the Crawler comprises two stages. In the first stage, we generate trajectories for a small subset of the training data and then perform imitation learning (see Appendix D.1 for details). In the second stage, reinforcement learning is applied. The details of the RL training implementation are described below.
@@
We conduct RL training on the AutoScholarQuery training set, where each instance consists of a query $q$ and a corresponding paper set $\mathcal{P}$. Starting with a query $q$, the Crawler generates a trajectory $\tau=(s_{1},a_{1},\cdots,s_{T},a_{T})$. At each time step $t$, we denote the current paper queue as $\mathcal{Q}_{t}$. Upon taking action $a_{t}$, the Crawler appends a set of new papers $(p_{1},p_{2},\cdots,p_{n_{t}})$ to the paper queue. If $a_{t}=\texttt{[Stop]}$, the set is empty and no papers are added.
@@
A key challenge in training the Crawler with RL is the significant time required to sample a complete trajectory for a given query. This is due to each [Search] or [Expand] action adding multiple papers to the paper queue, resulting in hundreds or even thousands of papers in the final paper queue.
To address this issue, we define a session as a sub-trajectory that ends with the [Stop] action, after which a new session begins. We identify two types of initial states for such sub-trajectories: $S_{q}$, containing only the user query, and $S_{q+p}$, containing both the query and a paper. $S_{q}$ represents the task’s starting point, where the LLM context includes only the query. In contrast, $S_{q+p}$ arises after a [Stop] action, where the LLM context is reset to the query and the next paper in the queue.
Formally, we model the Crawler as a policy $\pi_{\theta}(a_{t}|s_{t})$. We partition the entire trajectory $\tau=(s_{1},a_{1},\cdots,s_{T},a_{T})$ into a sequence of sessions: $(\tau_{t_{1}:t_{2}-1},\tau_{t_{2}:t_{3}-1},\cdots)$. Each session is $\tau_{t_{i}:t_{i+1}-1}=(s_{t_{i}},a_{t_{i}},\cdots,s_{t_{i+1}-1},a_{t_{i+1}-1})$, where the initial state $s_{t_{i}}$ is either belonging to type $S_{q}$ or $S_{q+p}$, and the final action $a_{t_{i+1}-1}$ is [STOP].
@@
$\hat{R}_{t}$ $=$ $\sum_{k=t}^{t_{i+1}-1}\gamma_{0}^{k-t}\bigg[r(s_{k},a_{k})$
$+\gamma_{1}\sum_{j=1}^{n_{k}}\hat{V}_{\phi}(S_{q+p_{j}})\bigg]-\beta\cdot\log\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\rm sft}(a_{t}|s_{t})}$
Here, $\gamma_{0}$ is the in-session discount factor, and $\gamma_{1}$ is the across-session discount factor. $\hat{V}_{\phi}(\cdot)$ is the value function model to approximate the state value. After executing $a_{k}$, the paper queue is updated to include the newly found papers $(p_{1},p_{2},\cdots,p_{n_{k}})$. Since the Crawler will subsequently initiate new sessions to process these additional papers, their associated reward-to-go should be incorporated into the return estimate. In addition, we include a per-token KL penalty term from the learned policy $\pi_{\theta}$ to the initial policy $\pi_{\rm sft}$ obtained through imitation learning at each token to mitigate over-optimization. This term is scaled by the coefficient $\beta$.
@@ 4.3 Selector
We perform imitation learning to optimize the Selector. See Appendix E for training data collection and training details.
@@ 5.1 Experimental Setting
The Selector was fine-tuned using the training dataset described in Appendix E. We conducted supervised fine-tuning for one epoch with a learning rate of 1e-5 and a batch size of 4. The training runs on 8 NVIDIA-H100 GPUs.
@@
The training process involves two stages. First, we perform imitation learning for 1 epoch on 12,989 training data with a learning rate of 1e-5 and batch size of 4 per device, using 8 NVIDIA H100 GPUs. In the second stage, we apply PPO training. To ensure stability, we first freeze the policy model and train the value model, followed by co-training both the policy and value models. The hyperparameters used during the training process are listed in Table 12 in Appendix D.2.
During imitation learning, the model encounters 5,000 queries, while during the RL training phase, the model processes a total of 16,000 queries. For more details please refer to Appendix D.1 for the imitation learning data construction and Appendix D.2 for the PPO training data sampling.
@@
The LLM predicts a query based on the context. Then the agent calls Google [^7] Accessed via the Google Search API provided by https://serper.dev. with the parameters site:arxiv.org and before:query_date, restricting search results by source and publication time.
@@
Table 4: Results on AutoScholarQuery test set. *: Due to the need for manual query submission, the ChatGPT baseline is evaluated on 100 randomly sampled instances. Results for all methods on this subset are reported in Table 14.
@@ 5.2 Baselines and Evaluation
  • ChatGPT. We submit scholar query to ChatGPT [^9] https://chatgpt.com, powered by search-enabled GPT-4o.
  • GPT-o1. Prompt GPT-o1 to process the scholar query. Note that it does not have access to external search tools.
  • PaSa-GPT-4o. Implement PaSa as illustrated in Figure 1 by prompting GPT-4o. It can perform multiple searches, paper reading, and citation network crawling.
We carefully designed prompts for all baselines and they are shown in Appendix H.2. All baselines, except PaSa-GPT-4o, represent the best-known scholar search methods. These comparisons highlight the effectiveness of our agentic approach. The comparison with PaSa-GPT-4o isolates the impact of RL training.
As shown in Figure 2, the crawling process of PaSa can be visualized as a paper tree. In practice, considering the computational expense, we limit the Crawler’s exploration depth to three for both PaSa-7b and PaSa-GPT-4o.
For Google-based baselines, we evaluate recall using Recall@20, Recall@50, and Recall@100 metrics for the top-20, top-50, and top-100 search results, respectively. For other baselines that do not produce rankings, we assess precision and recall for the final retrieved papers. Additionally, we compare Crawler recall between PaSa-GPT-4o and PaSa-7b, defined as the proportion of target papers collected by the Crawler. This measures how many target papers are successfully included in the paper queue generated by the Crawler.